# Search for echoes on the edge of quantum black holes

### v2

@@ Abstract
I Introduction
Post-merger gravitational wave (GW) echoes are our most direct observational windows into the quantum structure of black hole (BH) event horizons Cardoso et al. (2016a, b); Abedi et al. (2020), while their non-existence would rule out different hypotheses about the nature of these enigmatic objects. The best view of these horizons can be achieved by combining a large number of binary black hole (BBH) merger events. As such, the GW data release for BBH mergers during the first, second and third observing run of LIGO/Virgo observatories Abbott et al. (2019a, 2021a, 2021b, 2016a, 2016b, 2017, b) provides an unprecedented opportunity to test classical general relativity (GR), as well as its alternatives, in the strong gravity regime. One can assume GR as the base model and contrast it to GR+phenomenological echo waveforms to see which one fits the data better. Nonetheless, despite many attempts in searching for echoes Abedi et al. (2017a); Conklin et al. (2018); Abedi and Afshordi (2019); Uchikata et al. (2019); Holdom (2020); Westerweck et al. (2018); Nielsen et al. (2019); Salemi et al. (2019); Lo et al. (2019); Tsang et al. (2019); Abbott et al. (2020); Wang and Piao (2020); Westerweck et al. (2021); Ren and Wu (2021); Abbott et al. (2021c); Abedi et al. (2021); Conklin and Afshordi (2021) using different models, we still lack a waveform as physical/accurate as waveforms in GR. Additionally, there is no consensus on the optimal procedure to combine the events, for best sensitivity to fundamental physics. Although, the reported GW detections have so far been consistent with predictions of GR Abbott et al. (2016b, 2017, 2019b), the first search for echoes Abedi et al. (2017a), motivated by a resolution to the BH quantum information paradox, was conducted for the first observing run of the Advanced LIGO detectors (O1), which then motivated further searches within different GW data analysis frameworks, and using more physical echo waveforms. Accordingly, several attempts with replication and extention were made with positive Abedi et al. (2017a); Conklin et al. (2018); Abedi and Afshordi (2019); Uchikata et al. (2019); Holdom (2020); Conklin and Afshordi (2021), mixed Westerweck et al. (2018); Nielsen et al. (2019); Salemi et al. (2019), and negative Lo et al. (2019); Uchikata et al. (2019); Tsang et al. (2019); Abbott et al. (2020); Wang and Piao (2020); Miani et al. (2023); Uchikata et al. (2023) results. These searches lead to tentative evidence and detection found with different groups Abedi et al. (2017a); Conklin et al. (2018); Westerweck et al. (2018); Abedi and Afshordi (2019); Salemi et al. (2019); Uchikata et al. (2019); Holdom (2020); Abedi et al. (2021); Conklin and Afshordi (2021) at false alarm rates of $0.002\%-5\%$ (but see Westerweck et al. (2018); Ashton et al. (2016); Abedi et al. (2017b, 2018); Salemi et al. (2019); Abedi and Afshordi (2020); Abedi et al. (2020) for the ongoing discussion, comments, and rebuttals on statistical significance of these findings that motivate further investigations). So far, the searches for echoes have employed three strategies that can be classified into:
@@ I Introduction
    Waveform dependent Abedi et al. (2017a); Westerweck et al. (2018); Nielsen et al. (2019); Lo et al. (2019); Uchikata et al. (2019); Abbott et al. (2020); Wang and Piao (2020); Westerweck et al. (2021); Abbott et al. (2021c); Abedi et al. (2021); Uchikata et al. (2023).
@@
    Model-agnostic or coherent Abedi and Afshordi (2019); Conklin et al. (2018); Salemi et al. (2019); Holdom (2020); Ren and Wu (2021); Abedi et al. (2021); Conklin and Afshordi (2021); Miani et al. (2023).
@@
    Electromagnetic confirmation by Gill et al. Gill et al. (2019).
For more details, discussions and review please see Abedi et al. (2020).
Figure 1: Schematic diagram of GW echoes (stimulated Hawking radiation) from remnant of a BBH merger.
@@
A confirmed detection of echoes would imply that the BH horizon is not totally absorbing. This would lead to post-merger repeating signals which are produced in the cavity that traps GWs between the classical angular momentum barrier and the near-horizon membrane/firewall Abedi et al. (2017a); Cardoso et al. (2016a, b); Wang and Afshordi (2018); Abedi et al. (2021). However, firewalls are not a necessary condition to have observable echoes Abedi (2022). Stimulated emission of Hawking radiation, caused by the GWs that excite the quantum BH microstructure has a similar effect Oshita et al. (2020); Wang et al. (2019); Xin et al. (2021); Srivastava and Chen (2021); Abedi et al. (2021). As shown in Figs. 1 and 2, the trapped GWs slowly leak out, leading to repeating echoes within time intervals of:
@@
where $M_{\rm BH}$ and $a$ are the mass and the dimensionless spin of the final BH remnant. Here, $\ell_{\rm QG}$ is the characteristic physical length scale for quantum gravity effects where GWs are reflected near the (would-be) horizon. For $\ell_{\rm QG}=\ell_{\rm Planck}$, the reflection happens at a Planck distance from the horizon. More generally, for $\ell_{\rm QG}=\ell_{\rm Planck}/\Lambda$, deviations from GR happen sub-Planckian $\Lambda>1$ or super-Planckian $\Lambda<1$ scales. In this paper, we choose a conservative prior $-13\leq\log_{10}\Lambda\leq 13$.
Here, in comparison to former attempts, I used a more physical waveform, based on stimulated Hawking radiationOshita et al. (2020); Wang et al. (2019) to test for the existence of echoes. Furthermore, I adopt the Bayesian methodology and p and model, as in Abedi et al. Abedi et al. (2021).
In this approach, I set our model and search pipeline from a rather novel point of view to look for echoes combining 47 LVK BBH merger events, ensureing that these events meet the criterion, guaranteeing a false alarm rate $\rm{FAR}<10^{-3}\rm{yr}^{-1}$. I perform the search for echoes on BBH signals using the GWTC-1 Abbott et al. (2019a), GWTC-2 Abbott et al. (2021a) and GWTC-3 Abbott et al. (2021b). This search includes almost the bulk of all the confident observations Abbott et al. (2019a, 2021a, 2021b). The missed events are either the marginal ones or needed a high computational cost (ones with very small mass). In this approach of combining events I assume echo model is the same for all the events. In particular, I assume all the events have same echo amplitude $A$. Although, this approach does not cover entire space of former searches, it makes a complementary search in overall.
One such proposal to search for quantum black holes in GW data is given by phenomenological Boltzmann echoes waveform Oshita et al. (2020); Wang et al. (2019), where the general relativistic prediction for GW signal from BBH mergers $h_{\rm GR}(\omega)$ in Fourier space is modified to:
@@
where $Ae^{i\phi}$ quantifies their overall amplitude, while the modulus and phase of ${\cal R}$ quantify their relative damping and temporal separation, respectively. Generally, we expect $0\lesssim A\lesssim 2$ and $0<\phi\leq 2\pi$ due to GR non-linearities. Here $T_{H}$ is Hawking temperature of the black hole. Furthermore, I set the horizon mode frequency $m\times\Omega_{H}$ to m=2 for quadrupolar gravitational radiation (with the assumption that the energy in BBH ringdown and echoes are dominated by this mode) as main frequency of search pipeline. In this search we only keep n=2 number of echoes. Note that the time window to look for echoes is dictated by the model in (3), where its dependence on $\Delta t_{\rm{echo}}$ and mass dependence is given by (1). In this equation the final mass of black hole $M_{BH}$ is dictated by GR part of the event. The most physical and accurate term in this waveform is the Boltzmann factor ${\cal R}$ which resembles the same factor in Hawking radiation process. This term is responsible to stimulated Hawking radiation making this process observable by gravitational wave detectors. Indeed, the remaining assumptions made in the construction of this phenomenological waveform are less accurate because achieving greater precision necessitates numerical relativity computations.
Although, there is no doubt that the $h_{\rm GR}(\omega)$ (main event GR part) exists, we want to answer whether the echoes part exists. Existence of $h_{\rm GR}(\omega)$ helps us to obtain physical prior for echo model i.e. improvements in priors for $\Delta t_{\rm echo}$ in (1) or $\Omega_{H}$ and $T_{H}$ variables. Here, the Boltzmann factor $\exp[-\frac{|\omega-2\Omega_{H}|}{2T_{H}}]$ originates from Hawking tunnelling rate to fuzzy states of quantum BH (please see Fig 2 for this waveform). Note that, repeating echoes time delay modifies this factor to $\exp[\frac{\omega-2\Omega_{H}}{2T_{H}}+i\omega\Delta t_{\rm echo}]$ Oshita et al. (2020); Wang et al. (2019). Here, I only keep first two echoes of the waveform in this search pipeline. Indeed, this waveform is not as perfect as GR waveform, while it helps us in future research and establishment of better waveforms.
Data presented in this paper and configuration files to replicate the results (including priors for GR parameters) available at Sea .
Next section describes method and search pipeline and results. Section III provides an alternative approach to combine events hierarchically. Section IV, false/true positive estimations are provided for GW190521. Finally, I conclude with the search results and findings.
II Method and search pipeline and results
I employed PyCBC inference Biwer et al. (2019) pipeline using a dynamic nested sampling algorithm, dynesty [^1] I used 25,000 live points in each run. Speagle (2020). It is based on sampling the likelihood function for a hypothesis that gives a measure of existence of a signal in the data. The likelihood function that requires a power spectral density (PSD) is supposed to be compatible with the natural assumption that the background is Gaussian. For further details and the definition of the likelihood function in PyCBC Inference, please refer to Section 2.2 of Biwer et al. (2019). In order to combine the events, the amplitude A for all the 47 events (they are the events satisfy the criterion of a false alarm rate $\rm{FAR}<10^{-3}\rm{yr}^{-1}$) is fixed to a universal value and the individual Bayes factors of events BEvent(A) are combine as follows.
@@ II Method and search pipeline and results
[…]
I have used two/three detector networks H1-L1/H1-L1-V1 (Hanford, Livingston, Virgo) depending on the event and available data for this analyse GWT . While we know that LIGO/Virgo detectors are not stationary and/or Gaussian, as this was shown in Abedi et al. (2017a) even for short range of data, we assume that the PSD around each event to be our safest choice. In order get a more accurate estimation of the likelihood, the PSDs of each event are obtained from data around each of the events which is (-256,256) second interval centered on merger. Note that there are some events that this range cannot be applicable due to each detector operation. So, in order to obtain the PSD around these events, this range is slightly adjusted. First I obtain the Bayes factor comparing the log likelihood to the log likelihood of the Gaussian noise. Then the combined Bayes factors of alternative models with different amplitude A are compared (here they are $h_{\rm GR+echoes}(A,\omega)$ and $h_{\rm GR+echoes}(A=0,\omega)=h_{\rm GR}(\omega)$ in (2)). I used class of phenomenological IMR waveform family IMRPhenomPv2 Khan et al. (2016); Hannam et al. (2014) which is freely available as part of LALSuite LAL . Although the main search in Abedi et al. Abedi et al. (2021) for GW190521 has been performed with NRSur7dq4 waveform Varma et al. (2019), in order to make identical search with other events in this paper I employed IMRPhenomPv2 for this event. The slight change in reported Bayes factor for this event in this paper is due to change in waveform. It is worth to mention that other waveforms/changes have shown consistent result for this event Abedi et al. (2021). Note that the details on the results and smoke tests of the analysis for single event and its posterior for other parameters can be found in Abedi et al. (2021).
@@
[…]
GW150914 -0.53 GW170608 0.05 GW170818 -0.06
GW151226 -0.09 GW170809 0.08 GW170823 -0.25
GW170104 0.13 GW170814 -0.30
GW190408_181802190408_181802190408\_181802190408 _ 181802 -0.16 GW190521 0.96 GW190727_060333190727_060333190727\_060333190727 _ 060333 -0.30
GW190412 -0.09 GW190521_074359190521_074359190521\_074359190521 _ 074359 -0.54 GW190728_064510190728_064510190728\_064510190728 _ 064510 -0.01
GW190421_213856190421_213856190421\_213856190421 _ 213856 0.21 GW190602_175927190602_175927190602\_175927190602 _ 175927 -0.22 GW190814 -0.42
GW190503_185404190503_185404190503\_185404190503 _ 185404 -0.02 GW190630_185205190630_185205190630\_185205190630 _ 185205 -0.17 GW190828_063405190828_063405190828\_063405190828 _ 063405 0.04
GW190512_180714190512_180714190512\_180714190512 _ 180714 -0.06 GW190706_222641190706_222641190706\_222641190706 _ 222641 -0.06 GW190828_065509190828_065509190828\_065509190828 _ 065509 -0.14
GW190513_205428190513_205428190513\_205428190513 _ 205428 -0.15 GW190707_093326190707_093326190707\_093326190707 _ 093326 -0.02 GW190910_112807190910_112807190910\_112807190910 _ 112807 -0.30
GW190517_055101190517_055101190517\_055101190517 _ 055101 0.07 GW190708_232457190708_232457190708\_232457190708 _ 232457 -0.01 GW190915_235702190915_235702190915\_235702190915 _ 235702 -0.09
GW190519_153544190519_153544190519\_153544190519 _ 153544 -0.35 GW190720_000836190720_000836190720\_000836190720 _ 000836 -0.07 GW190924_021846190924_021846190924\_021846190924 _ 021846 0.00
GW191109_010717191109_010717191109\_010717191109 _ 010717 -0.36 GW191222_033537191222_033537191222\_033537191222 _ 033537 -0.32 GW200219_094415200219_094415200219\_094415200219 _ 094415 -0.07
GW191129_134029191129_134029191129\_134029191129 _ 134029 0.01 GW200112_155838200112_155838200112\_155838200112 _ 155838 -0.28 GW200224_222234200224_222234200224\_222234200224 _ 222234 -0.34
GW191204_171526191204_171526191204\_171526191204 _ 171526 0.01 GW200129_065458200129_065458200129\_065458200129 _ 065458 -0.43 GW200225_060421200225_060421200225\_060421200225 _ 060421 -0.01
GW191215_223052191215_223052191215\_223052191215 _ 223052 0.2 GW200202_154313200202_154313200202\_154313200202 _ 154313 0.21 GW200311_115853200311_115853200311\_$1$ 115853 -0.37
GW191216_213338191216_213338191216\_213338191216 _ 213338 0.03 GW200208_130117200208_130117200208\_130117200208 _ 130117 0.08 GW200316_215756200316_215756200316\_215756200316 _ 215756 -0.01
Table 1: Results of Bayes factor for GW echoes in GWTC-1, GWTC-2, and GWTC-3 events. Positive value of the $\log_{10}$ Bayes factor indicates a preference for the GR+echoes model over GR model, while the negative value suggests instead a preference for the GR model over the GR+echoes model. Here GW190521 shows loudest echo. Here based on Kass and Raftery (1995) all the individual events appear as inconclusive to both GR or GR+echoes with GW190521 as exception! (see Fig. 4).
Figure 4: Histogram of $\log_{10}$ Bayes factors of 47 events in Table 1. Vertical regions identify Jeffreys scale for interpretation of Bayes factor Kass and Raftery (1995).
@@
[…]
@@
The retrieved parameters and their associated errors align well with the reported values for each event GWOSC.
@@
Let us compare the Bayes factors in Table 1 with the common events analyzed using p-value estimation by the LIGO-Virgo-KAGRA (LVK) Collaboration Abbott et al. (2021c), as shown in Fig. 5. The plot in Fig. 5 also includes a degree 3 polynomial fit, suggesting a potential relationship between the LVK p-values and log10 Bayes factors obtained from gravitational wave echo analysis. The scatter of points around the fit highlights residual variability, indicating that while a potential trend exists, other factors such as noise, the inherent complexity of gravitational wave signals, and differences in models and methods may influence the correlation.
Figure 5: Comparison of LVK p-values Abbott et al. (2021c) and log10 Bayes factors in Table 1 for common events, with a degree 3 polynomial fit illustrating their gradual relationship in gravitational wave echo analysis. The curved trend suggests a potential correlation, indicating that higher Bayes factors may correspond to lower p-values, despite differences in methods and models.
III Combination of Bayes factors assuming a target distribution

### v3

@@ Abstract
1 Introduction
Post-merger gravitational wave (GW) echoes are our most direct observational windows into the quantum structure of black hole (BH) event horizons [1, 2, 3], while their non-existence would rule out different hypotheses about the nature of these enigmatic objects. The best view of these horizons can be achieved by combining a large number of binary black hole (BBH) merger events. As such, the GW data release for BBH mergers during the first, second and third observing run of LIGO/Virgo observatories [4, 5, 6, 7, 8, 9, 10] provides an unprecedented opportunity to test classical general relativity (GR), as well as its alternatives, in the strong gravity regime. One can assume GR as the base model and contrast it to GR+phenomenological echo waveforms to see which one fits the data better. Nonetheless, despite many attempts in searching for echoes [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] using different models, we still lack a waveform as physical/accurate as waveforms in GR. Additionally, there is no consensus on the optimal procedure to combine the events, for best sensitivity to fundamental physics. Although, the reported GW detections have so far been consistent with predictions of GR [8, 9, 10], the first search for echoes [11], motivated by a resolution to the BH quantum information paradox, was conducted for the first observing run of the Advanced LIGO detectors (O1), which then motivated further searches within different GW data analysis frameworks, and using more physical echo waveforms. Accordingly, several attempts with replication and extention were made with positive [11, 12, 13, 14, 15, 27], mixed [16, 17, 18], and negative [19, 14, 20, 21, 22, 28, 29] results. These searches lead to tentative evidence and detection found with different groups [11, 12, 16, 13, 18, 14, 15, 26, 27] at false alarm rates of $0.002\%-5\%$ (but see [16, 30, 31, 32, 18, 33, 3] for the ongoing discussion, comments, and rebuttals on statistical significance of these findings that motivate further investigations). So far, the searches for echoes have employed three strategies that can be classified into:
@@ 1 Introduction
    Waveform dependent [11, 16, 17, 19, 14, 21, 22, 23, 25, 26, 29].
@@
    Model-agnostic or coherent [13, 12, 18, 15, 24, 26, 27, 28].
@@
    Electromagnetic confirmation by Gill et al. [34].
For more details, discussions and review please see [3].
@@
A confirmed detection of echoes would imply that the BH horizon is not totally absorbing. This would lead to post-merger repeating signals which are produced in the cavity that traps GWs between the classical angular momentum barrier and the near-horizon membrane/firewall [11, 1, 2, 35, 26]. However, firewalls are not a necessary condition to have observable echoes [36]. Stimulated emission of Hawking radiation, caused by the GWs that excite the quantum BH microstructure has a similar effect [37, 38, 39, 40, 26]. As shown in Fig. 1, the trapped GWs slowly leak out, leading to repeating echoes within time intervals of:
@@
where $M_{\rm BH}$ and $a$ are the mass and the dimensionless spin of the final BH remnant. Here, $\ell_{\rm QG}$ is the characteristic physical length scale for quantum gravity effects where GWs are reflected near the (would-be) horizon. For $\ell_{\rm QG}=\ell_{\rm Planck}$, the reflection happens at a Planck distance from the horizon. More generally, for $\ell_{\rm QG}=\ell_{\rm Planck}/\Lambda$, deviations from GR happen sub-Planckian $\Lambda>1$ or super-Planckian $\Lambda<1$ scales.
Here, in comparison to former attempts, I used a more physical waveform, based on stimulated Hawking radiation[37, 38] to test for the existence of echoes. Furthermore, I adopt the Bayesian methodology and p and model, as in Abedi et al. [26].
In this approach of combining events I assume echo model is the same for all the events. In particular, I assume all the events have same echo amplitude $A$. Although, this approach does not cover entire space of former searches, it makes a complementary search in overall.
A promising approach for probing quantum‐modified horizons in gravitational‐wave data uses the phenomenological “Boltzmann echoes” waveform [37, 38]. Its central ingredient is the Boltzmann factor $\exp[-\frac{|\omega-2\Omega_{H}|}{2T_{H}}]$ which resembles the thermal weighting in Hawking’s radiation spectrum and governs the partial reflection of waves near the black‑hole horizon. This term is responsible to stimulated Hawking radiation making this process observable by gravitational wave detectors. Here, the Boltzmann factor originates from Hawking tunnelling rate to fuzzy states of quantum BH (please see Fig. 1 for this waveform). Note that, repeating echoes time delay modifies this factor to $\exp[\frac{\omega-2\Omega_{H}}{2T_{H}}+i\omega\Delta t_{\rm echo}]$ [37, 38].
Next section describes method and search pipeline and false/true positive estimations for GW190521. Section 3 provides combined Bayes factor estimation and an alternative approach to combine events hierarchically assuming a target distribution. Finally, I conclude with the search results and findings.
2 Method and search pipeline
Adopting a novel perspective, our model and search pipeline target echo signatures in 47 LVK BBH merger events, ensuring that these events meet the criterion, guaranteeing a false alarm rate $\rm{FAR}<10^{-3}\rm{yr}^{-1}$. I perform the search for echoes on BBH signals using the GWTC-1 [4], GWTC-2 [5] and GWTC-3 [6]. This search includes almost the bulk of all the confident observations [4, 5, 6]. The missed events are either the marginal ones or needed a high computational cost (ones with very small mass).
One such proposal to search for quantum black holes in GW data is given by phenomenological Boltzmann echoes waveform [37, 38], where the general relativistic prediction for GW signal from BBH mergers $h_{\rm GR}(\omega)$ in Fourier space is modified to:
@@ 2 Method and search pipeline
[…]
Although, there is no doubt that the $h_{\rm GR}(\omega)$ (main event GR part) exists, we want to answer whether the echoes part exists. Existence of $h_{\rm GR}(\omega)$ helps us to obtain physical prior for echo model i.e. improvements in priors for $\Delta t_{\rm echo}$ in (1) or $\Omega_{H}$ and $T_{H}$ variables.
I employed PyCBC inference [43] pipeline using a dynamic nested sampling algorithm, dynesty [^1] I used 25,000 live points in each run. [44]. It is based on sampling the likelihood function for a hypothesis that gives a measure of existence of a signal in the data. The likelihood function that requires a power spectral density (PSD) is supposed to be compatible with the natural assumption that the background is Gaussian. For further details and the definition of the likelihood function in PyCBC Inference, please refer to Section 2.2 of [43]. In order to combine the events, the amplitude A for all the 47 events (they are the events satisfy the criterion of a false alarm rate $\rm{FAR}<10^{-3}\rm{yr}^{-1}$) is fixed to a universal value and the individual Bayes factors of events BEvent(A) are combine as follows.
@@
I have used two/three detector networks H1-L1/H1-L1-V1 (Hanford, Livingston, Virgo) depending on the event and available data for this analyse [45]. While we know that LIGO/Virgo detectors are not stationary and/or Gaussian, as this was shown in [11] even for short range of data, we assume that the PSD around each event to be our safest choice. In order get a more accurate estimation of the likelihood, the PSDs of each event are obtained from data around each of the events which is (-256,256) second interval centered on merger. Note that there are some events that this range cannot be applicable due to each detector operation. So, in order to obtain the PSD around these events, this range is slightly adjusted. First I obtain the Bayes factor comparing the log likelihood to the log likelihood of the Gaussian noise. Then the combined Bayes factors of alternative models with different amplitude A are compared (here they are $h_{\rm GR+echoes}(A,\omega)$ and $h_{\rm GR+echoes}(A=0,\omega)=h_{\rm GR}(\omega)$ in (2)). I used class of phenomenological IMR waveform family IMRPhenomPv2 [46, 47] which is freely available as part of LALSuite [48]. Although the main search in Abedi et al. [26] for GW190521 has been performed with NRSur7dq4 waveform [49], in order to make identical search with other events in this paper I employed IMRPhenomPv2 for this event. The slight change in reported Bayes factor for this event in this paper is due to change in waveform. It is worth to mention that other waveforms/changes have shown consistent result for this event [26]. Note that the details on the results and smoke tests of the analysis for single event and its posterior for other parameters can be found in [26].
@@
    For each event listed in Table 1, calculate the Bayes factor by average integration over the amplitudes: $\mathcal{B}^{\rm{GR+echo}}_{\rm{GR}}=\frac{1}{A_{max}-A_{min}}\int_{A_{min}}^{A_{max}}B(A)dA$ with $(A_{min},A_{max})=(0,2)$. The result is presented in Table 1.
@@
Please find posterior distribution of General Relativity (GR) parameters at a minimum echo amplitude of $A=10^{-4}$ in supplemental material [41]. The retrieved parameters and their associated errors align well with the reported values for each event [45].
@@
In addition to the primary combined Bayes factor method (4), next section introduces an alternative hierarchical Bayesian approach. This method explicitly models the distribution of echo amplitudes across events using a target Gaussian distribution $\bar{\mathcal{P}}(\rm{A},\bar{A},\sigma)$. By hierarchically combining individual event Bayes factors while accounting for selection biases and parameter uncertainties, this approach infers the hyperparameters $\bar{A}$ (mean amplitude) and $\sigma$ (standard deviation) of the underlying population. Notably, this framework generalizes the primary method: the combined Bayes factor (4) emerges as the special case $\sigma\rightarrow 0$.
Data presented in this paper and configuration files to replicate the results (including priors for GR parameters) available at [42].
2.1 False positive and true positive estimation for GW190521
2.1.1 Using samples from GR posterior
This is the same analysis performed in [26]. In this section we perform the analysis using NRSur7dq4 waveform [49]. Thus, for consistency the Bayes factors of injections are compared to the one obtained with this waveform. Even though this section focuses on the outcome of the single event GW190521, it satisfies the validation test for ensuring the proper functioning of the method and model. Since in this part we don’t need to combine events and making a grid for amplitude A as the former sub-section, I made it as free parameter inside its prior range of the model $0\leq A\leq 2$. In order to evaluate the true positive we have to inject GR models containing echoes. The GR only model for injections are drawn from random samples given by GR posterior, which is obtained from 10 different seed runs. Note that in the GR run, due to the marginalization over the coalescence phase of the binary, this particular parameter was not included in the corresponding posterior sample. As a result, the random samples consist of 14 parameters, and the coalescence phase was estimated additionally using waveform fitting to the data. Once the GR part is selected out of random samples of its posterior, the echo part is added. The echo part is within a random selection from uniform priors of $13\leq\log_{10}\Lambda\leq 13$, $0\leq A\leq 2$, and $0\leq\phi\leq 2\pi$ which is given in previous part. In order to evaluate the background and minimize non-Gaussian and/or non-stationary effects the data around the event is being used. Therefore, they are injected at random offset times of $\pm[5,32]$sec compared to GW190521 merger time. The gap $\pm 5$sec around GW190521 is due to preventing any contamination of data. In order to compare the foreground (true positive) to the background (false positive), I also injected a sample with zero echo amplitude for any corresponding GR+echo injection at the same offset time. In the next step run the same script job for injections with GR+echo and GR waveform. The number of injections performed for zero amplitude echoes (GR waveforms) is 343, while this number is 110 for non-zero amplitude echoes (GR+echo waveform). Fig. 2 is the results of this analysis in a histogram plot (this figure is a re‐production of the data originally presented in Fig. 4 of [26]). We find injection Bayes factor for GR+echo with $34.5\pm 7.3\%$ and GR with $1.46^{+1.17}_{-0.88}\%$ being larger than $\mathcal{B}=7.5$ the echo Bayes factor for GW190521 with NRSur7dq4 waveform [26]. Then the empirical likelihood ratio $24.3^{+42.7}_{-11.8}$ which is the ratio of the foreground probability to the background probability is obtained.
Figure 2: Histograms to quantify false positive (GR injections) and true positive (GR+echo injections). Comparing to the GW190521 echo, we obtain their values as $1.46^{+1.17}_{-0.875}\%$, and $34.5\pm 7.3\%$ respectively. Note that this figure is a re‐production of the data originally presented in Fig. 4 of [26].
Uncertainties for these estimates (and throughout subsequent sections) are quantified via bootstrapping. To quantify sampling uncertainty in our true-positive and false-positive rates, we employ nonparametric bootstrap resampling of the injection ensembles. For each set of injections (main event and main event+echo), we generate 1,000 bootstrap replicates by resampling the original Bayes factor measurements with replacement. Within each replicate, we recompute the fraction of trials exceeding the evidence threshold $\log_{10}(B)=\log_{10}(7.5)$. The resulting bootstrap distributions provide uncertainty estimates, where we report the median along with 90% confidence intervals from the 5th–95th percentiles. Although uncertainties are substantial (e.g. $34.5\%\pm 7.3\%$ vs. $1.46^{+1.17}_{-0.88}\%$), the minimum true-positive rate ($\simeq 27\%$) remains an order of magnitude above the maximum false-positive rate ($\simeq 2.6\%$), confirming robust separation.
2.1.2 Using GW190521 signal
$\exp\left(-\frac{(t-t_{\mathrm{merger}})^{2}}{\xi^{2}}\right)\times\left(\begin{array}[]{c}\mathsf{Hanford\ data}\\ \mathsf{Livingston\ data}\\ \mathsf{Virgo\ data}\end{array}\right)$ (5)
where $t_{\rm{merger}}=1242442967.4$ sec (GPS time). The parameter $\xi$ should be taken in such that covers the main event signal and does’t leak to $\sim 1$ sec echo range. Thus $\xi=0.5$ sec is chosen. Then instead of GR part in former section we replace it with this extracted signal.
After preparing the injections, we follow the same procedure as described earlier, running the script for injections using both the GR+echo and GR waveforms. A total of 250 injections are performed for zero-amplitude echoes (main event injections), while 20 injections are conducted for non-zero amplitude echoes (main event+echo injections). The results of this analysis are presented in Fig. 3 as a histogram plot. We observe that the injection Bayes factor for the main event+echo injections and for main event injections being larger than $\mathcal{B}=7.5$ (echo Bayes factor for GW190521 with the NRSur7dq4 waveform [26]) is $35\pm 15\%$, and $4.4^{+2.0}_{-2.0}\%$ respectively. Subsequently, we calculate the empirical likelihood ratio as $7.8^{+7.8}_{-4.0}$, representing the ratio of foreground probability to background probability.
3 Results
The result for individual events and their histogram are reported in Table 1 and Fig. 4 respectively. Our measured Bayes factors for 46 events (out of 47 events) lie in the range 0.31.60.3–1.6, which falls into the “not worth more than a bare mention” regime ($10^{-0.5}\leq\mathcal{B}^{\rm{GR+echo}}_{\rm{GR}}\leq 10^{0.5}$) on Jeffreys scale [50] for interpretation of Bayes factor scale. Only one event GW190521 shows a Bayes factor of 9.2 of substantial evidence $10^{0.5}<\mathcal{B}^{\rm{GR+echo}}_{\rm{GR}}\leq 10$ on Jeffreys scale.
GW150914 -0.53 GW170608 0.05 GW170818 -0.06
GW151226 -0.09 GW170809 0.08 GW170823 -0.25
GW170104 0.13 GW170814 -0.30
GW190408_181802190408\_181802 -0.16 GW190521 0.96 GW190727_060333190727\_060333 -0.30
GW190412 -0.09 GW190521_074359190521\_074359 -0.54 GW190728_064510190728\_064510 -0.01
GW190421_213856190421\_213856 0.21 GW190602_175927190602\_175927 -0.22 GW190814 -0.42
GW190503_185404190503\_185404 -0.02 GW190630_185205190630\_185205 -0.17 GW190828_063405190828\_063405 0.04
GW190512_180714190512\_180714 -0.06 GW190706_222641190706\_222641 -0.06 GW190828_065509190828\_065509 -0.14
GW190513_205428190513\_205428 -0.15 GW190707_093326190707\_093326 -0.02 GW190910_112807190910\_112807 -0.30
GW190517_055101190517\_055101 0.07 GW190708_232457190708\_232457 -0.01 GW190915_235702190915\_235702 -0.09
GW190519_153544190519\_153544 -0.35 GW190720_000836190720\_000836 -0.07 GW190924_021846190924\_021846 0.00
GW191109_010717191109\_010717 -0.36 GW191222_033537191222\_033537 -0.32 GW200219_094415200219\_094415 -0.07
GW191129_134029191129\_134029 0.01 GW200112_155838200112\_155838 -0.28 GW200224_222234200224\_222234 -0.34
GW191204_171526191204\_171526 0.01 GW200129_065458200129\_065458 -0.43 GW200225_060421200225\_060421 -0.01
GW191215_223052191215\_223052 0.2 GW200202_154313200202\_154313 0.21 GW200311_115853200311\_115853 -0.37
GW191216_213338191216\_213338 0.03 GW200208_130117200208\_130117 0.08 GW200316_215756200316\_215756 -0.01
Table 1: Results of Bayes factor for GW echoes in GWTC-1, GWTC-2, and GWTC-3 events. Positive value of the $\log_{10}$ Bayes factor indicates a preference for the GR+echoes model over GR model, while the negative value suggests instead a preference for the GR model over the GR+echoes model. Here GW190521 shows loudest echo. Here based on [50] all the individual events appear as inconclusive to both GR or GR+echoes with GW190521 as exception! (see Fig. 4).
Figure 4: Histogram of $\log_{10}$ Bayes factors of 47 events in Table 1. Vertical regions identify Jeffreys scale for interpretation of Bayes factor [50].
This search differs from previous echo analyses in two key aspects: (1) the implementation of two combined methods, discussed below, and (2) the assumption of identical echo amplitudes across all gravitational wave events.
3.1 Hierarchical combination of Bayes factors
In order to obtain the overall Bayes factor of individual events and combined events we just do $\mathcal{B}^{\rm{GR+echo}}_{\rm{GR}}=\frac{1}{A_{max}-A_{min}}\int_{A_{min}}^{A_{max}}B(A)dA$ with $(A_{min},A_{max})=(0,2)$. For combined events I got $\mathcal{B}^{\rm{GR+echo}}_{\rm{GR}}\simeq 0.4$.
3.2 Hierarchical combination of Bayes factors assuming a target distribution
Figure 5: (a): Combined Bayes factor density in terms of amplitude for 47 events. Combined events give an overall value of $\mathcal{B}^{\rm{GR+echo}}_{\rm{GR}}\simeq 0.4$ for Bayes factor. (b): Combined Bayes factor posterior density where the Bayes factor curves are normalized by $\int_{A_{min}}^{A_{max}}B(A)dA$. Here, in order to compare the individual events (GW150914 and GW190521) and combined events we plot the posterior density of the Bayes factors.